OpenStudy (anonymous):

how to find vertical component of velocity if only given constant speed and time?

OpenStudy (stormfire1):

Well, unless the problem gives more information then you can't. You can however infer things from the wording of the problem. For example, if the problem is talking about a moving train, the vertical component of velocity would be 0 since the train isn't moving vertically.

OpenStudy (anonymous):

heres the problem, maybe im not seeing something : A crow is flying horizontally with a constant speed of 2.60 m/s when it releases a clam from its beak. The clam lands on the rocky beach 1.80 s later. Consider the moment just before the clam lands. (Neglect air resistance.) What is its vertical component of velocity?

OpenStudy (stormfire1):

Initially: Vx = 2.60 m/s Vy = 0 m/s acceleration (constant and down) = -9.8 m/s ^2 delta T = 1.80s For constant acceleration: Vf = Vi + at So for vertical directoin: Vf = 0 m/s -9.8 m/s * 2.80s Vf = -17.64 m/s (negative meaning a downward direction)

OpenStudy (stormfire1):

err...I typed 2.80s but meant 1.80s. However, the answer is still correct.

OpenStudy (anonymous):

awesome thanks a lot

OpenStudy (stormfire1):

No problem, good luck :)